OpenStudy (anonymous):

Two chords intersect inside a circle. The lengths of the segments of one chord are 4 and 6. The lengths of the segments of the other chord are 3 and __

OpenStudy (amistre64):

4*6 = 3*c

OpenStudy (amistre64):

what is 4*6?

OpenStudy (anonymous):

24

OpenStudy (amistre64):

24 = 3*c 24/3 = c

OpenStudy (anonymous):

8 = c

OpenStudy (amistre64):

then our missing value is 8

OpenStudy (anonymous):

Thanks!

OpenStudy (amistre64):

youre welcome, its just a little thrm i remember from geometry stuff .. one of the few :)